# Vamos: Versatile Action Models for Video Understanding

### v1

@@ 3.3 Lightweight Token Selector
Within the token selector, $\{x_{1}^{(i)},...,x_{n}^{(i)}\}$ are first projected to a lower dimension, then input into a shallow transformer encoder to obtain encodings $\{s_{1}^{(i)},...,s_{n}^{(i)}\}$. A linear layer then takes these encodings and generates the logits $g^{(i)}\in\mathbb{R}^{n}$ for final selection. During training, to maintain model differentiability, we apply Gumbel-Softmax [31] to the logits $g^{(i)}$ to pick the final selected token $x^{(i)}$ for each segment $X^{(i)}$. In this way, $k$ tokens $\{x^{(1)},...,x^{(k)}\}$ are sampled as a condensed representation of the original text sequence with much smaller length and input to downstream LLMs for video understanding.
@@ 4 Experiments
Table 1: Vamos with various video representations on LTA task. Results are reported on Ego4D LTA test set, the metric is edit distance.
Input Ego4D-LTA normal-$\downarrow$
Verb Noun Action
vision 0.653 0.673 0.884
action 0.661 0.651 0.878
caption 0.676 0.675 0.890
vis+act 0.643 0.650 0.868
Input NeXT-QA normal-$\uparrow$ IntentQA normal-$\uparrow$
Cau. Tem. Des. All CW CH TP&TN All
vision 61.5 58.2 66.3 61.2 55.2 62.4 46.9 54.4
action 55.1 53.4 53.2 54.3 57.5 65.2 47.4 56.3
caption 70.9 69.3 77.7 71.5 69.7 68.8 61.7 67.6
vis+cap 70.9 67.2 77.7 70.8 69.0 69.9 63.2 67.7
Table 1: Vamos with various video representations on LTA task. Results are reported on Ego4D LTA test set, the metric is edit distance.
Table 2: Vamos with various video representations on VQA task. Results are reported on NeXT-QA and IntentQA test set, the metric is accuracy. On both dataset, caption-based representations achieve competitive performance.
@@ 4.1 Task and Datasets
We focus on two major video understanding tasks: long-term action anticipation and video question answering. Long-term action anticipation. The long-term action anticipation (LTA) task asks a model to predict a sequence of actions in form of a verb-noun pairs in a long future window based on video observations of the past actions. In LTA, a long video $V$ is first split into a number of annotated video segments. Given video observation before segment $i$, our task is to predict the future actions in sequences of verb-noun pairs of the next $Z$ segments allowing $K$ candidate sequences. The correctness of the predicted sequence is measured with edit distance. We evaluation on:
@@
EgoSchema [50] is a recently proposed diagnostic datasets derived from Ego4D for long-form video understanding. In contrast to previous video question answering benchmarks where the input videos are generally less than one-minute long, Egoschema contains videos that span 3-minute long with high temporal certificate to testify the video understanding model’s capability to understand long videos. It contains 5031 videos and each video has a multiple choice question. It requires zero-shot evaluation and releases answers to 500 questions for development.
NeXT-QA [71] is a popular multiple choice video question answering benchmark that test video understanding in terms of describing and reasoning the temporal actions. It contains 5,440 video clips and 47,692 questions, grouped into causal (48%), temporal (29%) and descriptive (23%).
IntentQA [43] is another multiple choice VQA dataset built on top of NeXT-QA but focuses on intent reasoning. The authors select the videos related to causal and temporal questions from NeXT-QA and constructed their own questions and answers to focus on testifying models’s performance on reasoning questions.
@@ 4.2 Implementation
Generating Video Captions. We zero-shot generate task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema, LLaVA-1.5 for NeXT-QA. For Ego4D LTA, we sample the center frame for each video segment to generate its caption. The captions for the 8 observed segments are then concatenated as the video representations. For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, then caption each frame and concatenate these frame-level captions to represent the video. For NeXT-QA, IntentQA, and EgoSchema we sample 6, 6, and 12 frames respectively.
Vamos for Temporal Modeling. For full-shot VQA and LTA, We use Llama2-7B [64], an open-sourced LLM as the temporal model for video understanding. During training, we use low-rank adaption [26] (LoRA) adapter to perform parameter-efficient fine-tuning on the training set. For vision input, we use the frozen CLIP [58] ViT-L/14 to extract image features. For the zero-shot long-form VQA on EgoSchema, we use several popular LLMs including OpenAI GPT-3.5-turbo, GPT-4 and Llama2-chat-13B.
4.3 LTA: Action Modeling with Vamos
We first apply Vamos on the long-term action anticipation task, which requires direct modeling of video temporal dynamics by predicting future actions tokens based on video observation. We finetune Vamos on Ego4D LTA dataset with three types of representations: the recognized action labels, free-formed captions, and the continuous vision features. As detailed in the first three rows in Table 2, we observe that the action-based representation typically outperforms the vision-based input. Notably, the action representation, despite its compact nature, surpasses captions, which provide more extensive details and descriptions of the observed events. This suggests that a discrete and structured sequence of actions serves as an effective representation for modeling video temporal dynamics.
@@ 4.4 LLM as Long-term Video Temporal Reasoner
Our experiments on LTA task reveal that compact action labels are effective for video temporal modeling, but is action representation alone sufficient for comprehensive video understanding? We then delve into a more challenging task: long-form video question answering, which demands an extensive understanding and reasoning of the video content. We conduct zero-shot experiment on EgoSchema, a recent challenging video question answering benchmark where existing state-of-the-art visual-language models struggle.
Model Input Type Full Set Acc.
@@
Table 3: zero-shot VQA on Egoschema. * denotes subset results.
We employ OpenAI GPT-4 as the zero-shot video reasoner, and the same models used for LTA to zero-shot extract action labels and captions (BLIP-2) as well. Results are listed in Table 3. We observe the LLM utilizing both zero-shot generated actions and captions largely outperform current best-performed visual-language model (InternVideo [68]). Contrasting with the action prediction task, captions with abundant details and semantic information serve as a much stronger representation for the challenging VQA task. Furthermore, we use the video narrations of 50 examples provided by the authors as “oracle” captions and evaluate on this subset. Remarkably, the LLM achieves an impressive accuracy of 81.8% with the “oracle” captions. Although not directly comparable with the full-set performance, it reflects the vast potential and generalizability of caption-based textual video representations and the LLMs’ capability for reasoning over broad time spans.
@@ 4.5 What Makes Good Video Representation?
In addition to text features, Vamos can also integrate vision-language features. We then investigate whether different modalities encode complementary information on Ego4D, IntentQA, and Next-QA. We train the Vamos model to incorporate the visual embeddings (the linear layer in Figure 2), and use LoRA to update the weights of Llama-2.
From Ego4D LTA results shown in Table 2, we observe incorporating vision input with the compact action sequence further enhances model performance especially on verb and action prediction. For the VQA benchmarks NeXT-QA and IntentQA, as illustrated in Table 2, the caption-based textual video representations significantly outperform purely visual representations in both content description and causal reasoning. This aligns with our prior observations from EgoSchema that captions can serve as a strong representation for question answering and reasoning. Contrasting with the LTA task, adding visual features to captions only marginally influence the performance on the two VQA datasets. This suggests that the task-agnostic captions already contain sufficient information and cues to address the questions in these benchmarks.
4.6 How many tokens are needed for reasoning?
While we show zero-shot captions generated by open-box models are effective representations of video and offer a competitive interface to interact with LLMs, we further ask the following questions: How many tokens are really necessary to represent videos? What specific information from lengthy captions is actually being utilized by LLMs? We then implant a lightweight token selector to Vamos to automatically compress the captions into much more compact token sequences.
@@ 4.6 How many tokens are needed for reasoning?
We conduct experiment on the NeXT-QA benchmark. In detail, after the word embedding layer in Vamos, the long caption-based textual video token sequence is uniformly partitioned into 40 segments. The token selector are then applied on each segment to select one token. In this way, we condense a long sequence (644 on average) into 40 tokens. These 40 tokens are then fed into LLMs as the new compressed representation of original video for down-stream video question answering task. During training, the token selector is optimized together with the Vamos.
For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA. For comparison, we also show the vision-only performance taking in 12 frames and the performance with unselected caption-based input. We show the results in Table 4. After condensing the prompts from an average of 644 tokens to 40 tokens, Vamos experiences an expected performance drop but still achieves a competitive 65.3% accuracy, significantly outperforming the vision-only baseline utilizing double input frames by 4.1%. When further compressed to 20 tokens, the performance drops to 59.5%, slightly worse than the vision baseline. This shows the sparsity of the free-form natural language captions for video reasoning and indicates a promising potential to compress the textual representation for better practical application. Besides, instead of using hard selection (choosing one token with highest score) during inference, we also perform soft-selection (using weighted pooling by applying Softmax instead of hard-max to the logits). Interestingly, the soft-selection achieves similar performance with hard-selection even with access to the whole sequence, indicating there is few missing information left in the unselected tokens.
 Input selected # Cau. Tem. Des. All
Frame all 61.5 58.2 66.3 61.2
Caption all 70.9 69.3 77.7 71.5
Caption 20 / 644 59.5 56.8 64.6 59.5
Caption 40 / 644 65.4 62.4 70.5 65.3
Caption* 40 / 644 65.6 62.4 70.4 65.4
Table 4: Condensing captions with token selector on NeXT-QA. Average caption length is 644. * denotes soft-selection results.
Figure 3: Illustration of Vamos predictions and manual intervention.
Figure 4: Illustration of predictions with token selector and manual intervention. The selector picks tokens highly related to the question and intervention corrects wrong predictions without training.
4.7 Qualitative Analysis and Intervention
Compared with continuous visual features, one major advantage of the textual video representation is it’s excellent interpretability that making understanding the relationship between input and prediction direct and simple. We then make qualitative analysis of some positive and negative results from Vamos and the token selector.
@@ 4.7 Qualitative Analysis and Intervention
Visualization of results. In Figure 3 we demonstrate two correctly-answered examples and one incorrect of Vamos from IntentQA. From the two positive examples we can see that the generated captions well describe the scene and activities happening in the video (“overlooking water” and “playing keyboards”), thus providing strong clues for LLMs to answer the question about description (“look at scenery”) and reasoning (“play guitar with man playing keytar”). However, a suboptimal caption will also cause the Vamos’ failure of reasoning. In the third negative example, the captions successfully describe the baby’s action of eating but fail to acknowledge the presence and potential actions of a woman in the corner of the scene. This omission leads to an incorrect prediction regarding the woman’s reaction.
Visualization of Token Selector. By condensing the captions consists of thousands of tokens into 40 with a token selector, it not only significantly accelerates the training and inference speed, but also provides a better view for us to understand what’s the key element for Vamos to understand and reason about the video. In Figure 4, we demonstrate two positive examples and one negative example of Vamos with a token selector from NeXT-QA. After the token selector, captions in natural language are broken into separated words that highly descriptive and related to the question. The two positive examples demonstrate Vamos’ strong ability to reason by the selected tokens providing direct or indirect clues to the question. In the challenging negative example, even highly related tokens such as “race” and “track” are selected, Vamos still tends to relate “tyres” with the more common word “car”, indicating the potential bias encoded by the LLMs from pre-training.
@@
Test Time Intervention. In addition to good interpretability, the text-based white-box representation also provides the capability of post-hoc test-time intervention, which is pivotal for diagnosing and fixing failed predictions without retraining the reasoning model. In Figure 3 and Figure 4, we show examples of test-time intervention of the negative examples for both original Vamos and Vamos with selector. By providing more accurate and related captions or tokens, Vamos is able to correct the wrong answers.
@@ 4.8 Design Choices and Ablation Study
Caption models. We study the impact of caption models on the reasoning performance on NeXT-QA. We compare two captioning models: BLIP-2 and LLaVA-1.5-7b. We observe that captions generated by BLIP-2 are generally concise spanning less than 20 tokens long while their counterparts from LLaVA-1.5 are more detailed and descriptive with about 100 tokens on average. Results shown in Table 5 shows that captions from LLaVA-1.5 achieve better performance. Interestingly, the two captions achieve similar performance on descriptive questions but LLaVA-1.5 captions perform significantly better on casual and temporal reasoning. This indicates that reasoning from videos generally benefit more from extra details compared with describing and summarizing.
Number of Captions. We study the impact of sampled frame numbers for captioning on NeXT-QA. As shown in Table 5, we found that with more captions of frames spanning across the video the model will perform better. This result is align with our expectation as we bring information with more granularity from the videos.
Comparison of LLMs on zero-shot VQA Besides, we also compare various LLMs’ zero-shot performance based on the same caption representations on EgoSchema. As shown in Table 6, GPT-4 achieves huge improvements comparing with GPT-3.5-turbo and Llama2-Chat-7B, demonstrating the potentiality and advantage of textual representation to naturally incorporate the improvements of LLMs.
 Caption Frame # Cau. Tem. Des. All
LLaVA-1.5 1 66.4 63.0 72.0 66.3
LLaVA-1.5 3 68.5 65.2 76.4 68.8
LLaVA-1.5 6 70.9 69.3 77.7 71.5
BLIP-2 6 67.9 64.6 77.1 68.4
Table 5: Ablation on captioning model and frame numbers.

### v2

@@ 3.3 Token Bottleneck Models
Within the token selector, $\{z_{1}^{(i)},...,z_{n}^{(i)}\}$ are first projected to a lower dimension, and then provided as inputs to a shallow transformer encoder to obtain encodings $\{s_{1}^{(i)},...,s_{n}^{(i)}\}$. A linear layer then takes these encodings and generates the logits $\mathbf{g}^{(i)}\in\mathbb{R}^{n}$ for final selection. During training, we apply Gumbel-Softmax [28] on the logits $\mathbf{g}^{(i)}$ to pick the token $z^{(i)}$ for each segment $\mathbf{z}_{\text{tvr}}^{(i)}$ while ensuring the module is differentiable. In this way, $k$ tokens are selected as the condensed representation of the original tokenized input sequence $\mathbf{z}_{\text{tvr}}$.
The token selector in TBM allows us to inspect the important evidence selected for the downstream tasks, and to intervene the wrongly recognized or selected tokens with the correct ones with human in the loop (Figure 2 right). Practically, the token selector also can also speed up the inference time due to its own light-weight implementation, and that only a much smaller subset of the tokens (e.g. 6%) are processed by the computationally heavy LLM.
@@ 4.1 Task and Datasets
Long-term action anticipation. The LTA task asks a model to predict a sequence of actions in form of a verb-noun pairs in a long future window based on video observations of the past actions. In LTA, a long video $V$ is first split into a number of annotated video segments. Given video observation before segment $i$, our task is to predict the future actions in sequences of verb-noun pairs of the next $Z$ segments allowing $K$ candidate sequences. The correctness of the predicted sequence is measured with edit distance. We evaluation on:
@@
EgoSchema [50] is a recently proposed diagnostic datasets derived from Ego4D for long-form video understanding. Egoschema contains videos that span 3-minute long with high temporal certificate to testify the video understanding model’s capability to understand long videos. It contains 5031 videos and each video has a multiple choice question. It requires zero-shot evaluation and releases answers to 500 questions for development.
NeXT-QA [75] is a popular multiple choice video question answering benchmark that tests video understanding in terms of describing and reasoning the temporal actions. It contains 5,440 video clips and 47,692 questions, grouped into causal (48%), temporal (29%) and descriptive (23%).
IntentQA [39] is a multiple choice VQA dataset built on top of NeXT-QA but focuses on intent reasoning. The authors select the videos related to causal and temporal questions from NeXT-QA and constructed their own questions and answers to focus on testifying models’s performance on reasoning questions.
@@ 4.2 Implementation
Generating Video Captions. We generate zero-shot, task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema. We use LLaVA-1.5 to generate captions for NeXT-QA due to its better performance (see Table 4.5 for an ablation study). For Ego4D LTA, we sample the center frame for each video segment to generate its caption. The captions for the 8 observed segments are then concatenated as the video representations. For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, and then generate and concatenate the frame-level captions. We sample 6, 6, and 12 frames for NeXT-QA, IntentQA, and EgoSchema, respectively.
Vamos for Temporal Modeling. For full-shot VQA and LTA, We use LLaMA-7B and LLaMA2-7B [67] respectively as the temporal model for video understanding. During training, we use LLaMA Adapter [89] or low-rank adaption [23] (LoRA) to perform parameter-efficient fine-tuning on the training set. For vision input, we use the frozen CLIP [59] ViT-L/14 to extract image features. For the zero-shot long-form VQA on EgoSchema, we use several popular LLMs including OpenAI GPT-3.5-turbo, GPT-4, GPT-4o and LLaMA2-chat-13B.
Table 1: Vamos with various video representations. Results are reported on Ego4D LTA test set, the metric is edit distance. Results are reported on NeXT-QA validation set, and IntentQA test set, the evaluation metric is accuracy. On all datasets, text-based representations achieve competitive performance.
Input Ego4D-LTA $\downarrow$ NeXT-QA $\uparrow$ IntentQA $\uparrow$
Verb Noun Action Cau. Tem. Des. All CW CH TP&TN All
vision 0.653 0.673 0.884 69.6 67.2 74.7 69.6 68.9 71.6 58.0 66.7
text 0.661 0.651 0.878 75.5 71.3 81.1 75.0 74.0 78.6 67.5 73.2
vis+text 0.643 0.650 0.868 74.5 71.0 81.7 74.5 73.5 76.6 64.3 71.7
@@ 4.3 LLM as Long-term Video Temporal Reasoner
We first apply Vamos on the long-term action anticipation task, which requires direct modeling of video temporal dynamics by predicting future actions tokens based on video observation. We fine-tune Vamos on Ego4D LTA dataset with continuous visual embeddings or text-based representation. We observe that action-based representation slightly outperforms the free-form captions due to the nature of the task. As shown in the first three columns of Table 4.2, we observe that the text-based representation outperforms the vision-based input, and combining the two further improves the performance.
Table 2: zero-shot VQA on Egoschema. *: 500 question subset.
Model Input Type   Acc.
@@
Table 3: Ablation on the number of frames on EgoSchema with GPT-3.5 turbo.
# Frames   Full Set Acc.
1 37.83%
4 38.36%
12 41.24%
We now consider the more challenging task, long-form video question answering, by conducting zero-shot experiment on the recently collected EgoSchema benchmark. We employ OpenAI GPT-4 as the video “reasoner”. We observe that free-form captions extracted by BLIP-2 consistently outperform the action-based representation, presumably because richer information is retained in the captions. We report the free-form caption-based performance from now on, unless otherwise mentioned. From Table 4.3, we observe that Vamos with text-based video representation largely outperforms the state-of-the-art InternVideo model [71], which is jointly trained on vision and language inputs. We attribute the performance gain due to the decoupling of perception and reasoning, the latter of which we hypothesize is easier to generalize even at zero-shot, thanks to the LLM pre-training. To better understand the performance upper-bound of the text-based representation, we use 500 ground-truth video narrations provided by the authors of EgoSchema to perform evaluation on this subset. Remarkably, the LLM achieves an impressive accuracy of 81.8% with the oracle captions. Although not directly comparable with the full-set performance, the result confirms the potential of the Vamos framework for reasoning over broad time spans in the long-form video question answering task, and that better empirical performance may be achieved by improving the captioning models.
@@ 4.4 What Makes Good Video Representation?
In addition to text features, Vamos can also integrate vision-language features. We then investigate whether different modalities encode complementary information on Ego4D, IntentQA, and NeXT-QA. We perform parameter efficient fine-tuning to update the weights of the LLM, whose gradients are used to jointly train the linear projection layer to incorporate the visual embeddings. We observe that naively fine-tuning with text and visual inputs lead to model overfitting, and hence perform modality dropout (i.e. randomly discard the entire sequence of visual embeddings) when fine-tuning the vis+text models.
While visual and text inputs are complementary in the Ego4D LTA task, we observe in Table 4.2 that the caption-based representation significantly outperform the vision-only baseline in the NeXT-QA and IntentQA benchmarks for video question answering, and that adding visual features only marginally affects the performance. This suggests that pre-trained visual embeddings such as CLIP may not encode residual information useful for the video QA task (e.g. fine-grained visual and motion features). Extracting visual representations that would complement the information encoded by the task-agnostic captions is thus an important future work.
Figure 3: Visualizations of example Vamos predictions and manual intervention.
4.5 Token Bottleneck Models
Although the text-based video representation is directly interpretable, it remains unclear which tokens are selected as evidence for making predictions. The token bottleneck models (TBMs) enable us to reveal the selected evidence and also improve the model’s inference speed.
@@ 4.5 Token Bottleneck Models
VQA Performance. We first tokenize and embed the input sequence $\mathbf{x}_{t}$ to obtain $\mathbf{z}_{t}$, which is uniformly partitioned into $k=$ 20 or 40 segments. TBM is then applied on each segment to select one token. In this way, we condense a long sequence (644 tokens on average for NeXT-QA) into $k$ tokens. TBM is jointly optimized with Vamos during training.
For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA-1.5. For comparison, we also show the vision-only performance taking in 12 frames and the performance with unselected caption-based input. Table 4.5 shows that the TBM leads to an expected performance drop due to discarding over 90% of the input tokens, and increasing $k$ improves the TBM performance. When $k=40$, Vamos achieves a competitive 65.3% accuracy, while only adding 0.7M parameters to the Vamos framework and achieving 5x inference speedup from 1.41s to 0.29s per sample on a single A6000 GPU.
Figure 4: Illustration of the token bottleneck model and the impact of intervention on model’s predictions. The selected tokens are highly related to the task (examples $a$ and $b$) and can be manually intervened to correct the model’s prediction (example $c$).
Table 4: Condensing captions with token bottleneck models on NeXT-QA.
Input Selected tokens ($k$) Cau. Tem. Des. All
vision all 69.6 67.2 74.7 69.6
text all 75.5 71.3 81.1 75.0
text 20 / 644 59.5 56.8 64.6 59.5
text 40 / 644 65.4 62.4 70.5 65.3
@@
Visualization of Vamos predictions. Figure 3 provides two positive and one negative examples selected from IntentQA. The predictions are made by Vamos without TBM. From the two positive examples we can see that the generated captions manage to describe the scene and activities happening in the video (“overlooking water” and “playing keyboards”), thus providing strong clues for LLMs to answer the question about description (“look at scenery”) and reasoning (“play guitar with man playing keytar”). However, a sub-optimal caption will also cause the Vamos’s failure of reasoning. In the third example, we observe that the captions successfully describe the baby’s action of eating but fail to describe the presence and potential actions of a woman in the corner of the scene. This omission leads to an incorrect prediction regarding the woman’s reaction.
Visualization of the Selected Tokens. TBM not only helps accelerate Vamos’ inference speed, but also allows us to interpret which pieces of evidence are used by the model to make predictions. In Figure 4, we show two positive examples and one negative example of Vamos with TBM from NeXT-QA. The two positive examples demonstrate Vamos’ ability to select tokens that serve as evidence for question answering. The negative example is challenging: even highly related tokens such as “race” and “track” are selected by Vamos, it still fails to reason the purposes of the tyres in the video.
@@
Test Time Intervention. Text-based video representation is not only directly interpretable as used by Vamos and TBM, it also allows the users to conduct post-hoc, test-time intervention, which is pivotal for diagnosing and fixing failed predictions without retraining. In Figure 3 and Figure 4, we show examples of test-time intervention performed on the negative examples for Vamos and Vamos with TBM, respectively. By providing more accurate and related captions or tokens, Vamos is able to correct its wrong predictions.
Table 5: Ablation on captioning model and frame numbers on NeXT-QA.
 Caption # Frame Cau. Tem. Des. All
LLaVA-13B 1 69.9 67.2 74.8 69.8
LLaVA-13B 3 73.0 70.4 79.4 73.1
LLaVA-13B 6 75.5 71.3 81.1 75.0
LLaVA-7B 6 75.2 71.9 81.6 75.1
BLIP-2 6 72.7 68.9 78.8 72.4
Table 6: Comparison of different LLaMA models on NeXT-QA.
Model Cau. Tem. Des. All
LLaMA1-7B 75.5 71.3 81.1 75.0
LLaMA2-7B 74.8 72.3 81.6 75.0
LLaMA3-8B 77.2 75.3 81.7 77.3
@@ 4.6 Design Choices and Ablation Study
Caption Models. We study the impact of caption models on the video QA performance on NeXT-QA. We compare two captioning models: BLIP-2 and LLaVA-1.5. We observe that captions generated by BLIP-2 are generally more concise (less than 20 tokens), while captions generated by LLaVA-1.5 are more detailed (around 100 tokens on average). Results shown in Table 4.5 shows that captions from LLaVA-1.5 achieve better performance. We also investigate the influence of caption model size by comparing LLaVA-1.5 7B and 13B versions. Interestingly, scaling LLaVA from 7B to 13B does not lead to improvement.
Number of Frames. We study the impact of sampled frame numbers for captioning on NeXT-QA. As shown in Table 4.5, we found that using more captioned frames leads to better performance. A similar trend can be observed on EgoSchema in Table 4.3. However, we observe diminishing return when 12 frames are used, and hence not worth the speed and accuracy trade-off.
Impact of LLMs on VideoQA. We study the impact of different LLaMA versions on NeXT-QA. As shown in Table 4.5, Vamos directly benefits from a more advanced LLM, which is a desirable property for practitioners.
Impact of LLMs on Long-form VideoQA. We study the impact of LLMs on the EgoSchema benchmark. As shown in Table 4.7, GPT-4o achieves significant improvements comparing with GPT-4 and LLaMA2-Chat-7B baselines. This again demonstrates that Vamos can directly benefit from advances in the large language models, making its decoupled perception and reasoning design desirable for practitioners.